OpenStudy (anonymous):

2x+2y-z=9 3x-y+3z=9 x-2y+z=3 I know that x=1, but I don't know the rest.

OpenStudy (anonymous):

(2x-2y+z)+ (2x+2y-z)=3+9=12=4x x=3